# VALOR-Eval: Holistic Coverage and Faithfulness Evaluation of Large Vision-Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) suffer from hallucination issues, wherein the models generate plausible-sounding but factually incorrect outputs, undermining their reliability. A comprehensive quantitative evaluation is necessary to identify and understand the extent of hallucinations in these models. However, existing benchmarks are often limited in scope, focusing mainly on object hallucinations. Furthermore, current evaluation methods struggle to effectively address the subtle semantic distinctions between model outputs and reference data, as well as the balance between hallucination and informativeness. To address these issues, we introduce a multi-dimensional benchmark covering objects, attributes, and relations, with challenging images selected based on associative biases. Moreover, we propose an large language model (LLM)-based two-stage evaluation framework that generalizes the popular CHAIR metric Rohrbach et al. (2018) and incorporates both faithfulness and coverage into the evaluation. Experiments on 10 established LVLMs demonstrate that our evaluation metric is more comprehensive and better correlated with humans than existing work when evaluating on our challenging human annotated benchmark dataset. Our work also highlights the critical balance between faithfulness and coverage of model outputs, and encourages future works to address hallucinations in LVLMs while keeping their outputs informative.111Our dataset and code can be found here: https://github.com/haoyiq114/VALOR.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Liu et al. (2023b); OpenAI (2023) have shown remarkable performance across a broad range of vision-language tasks. Despite the promising progress, the issue of hallucinations has emerged as a critical concern. Hallucination refers to the generation of plausible sounding but inaccurate or fabricated textual descriptions for a given image, which can compromise the reliability and trustworthiness of the models.
@@
Recent studies have proposed various methods to evaluate models’ generative hallucinations Wang et al. (2023a); Zhai et al. (2023); Jing et al. (2023) and discriminative hallucinations Li et al. (2023b); Guan et al. (2023); Lovenia et al. (2023). However, they predominantly focus on hallucinations concerning object existence and their faithfulness within generated content, often neglecting other critical types of hallucinations and the assessment of coverage. This oversight can result in a lack of attention to the variety and depth of hallucinations that may occur beyond object identification, such as attributes and relations. Furthermore, these evaluation methods are often constrained by a predefined vocabulary, thus are inherently limited to fully appreciate the richness of the free-form generated captions. Specifically, the evaluation metrics may not capture novel expressions that extend beyond the predetermined vocabulary.
In contrast to prior studies, we introduce a human-annotated multi-dimensional evaluation benchmark VALOR-Bench222VALOR is short for vision-language attribute, relation, and object coverage and faithfulness evaluation. by breaking down hallucinations into three categories: object (existence), attributes (color and count), and relations (positional and comparative). In addition, to make the test cases challenging, we utilize the associative biases presented in training datasets to select images with only one component of commonly co-occurred pairs or groups, leading models to mistakenly generate associated elements that are not present. Our experimental findings validate the effectiveness of this methodology in exposing the susceptibility of current LVLMs to such biases.
@@
We evaluate 10 mainstream LVLMs on VALOR-Bench, focusing on the balance between faithfulness and coverage score. We notice that even GPT-4V(ision) OpenAI (2023) still suffers from hallucination, achieving a relatively low faithfulness score despite covering more information within an image comparing to other models.
@@ 2 Existing LVLMs Hallucination Evaluation Benchmarks and Metrics
As shown in Table 1, existing studies Li et al. (2023b); Wang et al. (2023a); Zhai et al. (2023); Lovenia et al. (2023); Villa et al. (2023) have primarily focused on object-level hallucination, with only a few recent studies Jing et al. (2023); Wang et al. (2023b) recognizing the importance of extending hallucinations to other dimensions. Our benchmark VALOR-Bench covers hallucination evaluations of objects, attributes, and relations, and we further detail attributes to color and counting, and relations to positional and comparative, to provide a comprehensive and fine-grained evaluation benchmark.
@@ 3.1 Image Collection
We aim to select images that can effectively expose the issue of model hallucinations. We hypothesize that when models are repeatedly exposed to specific combinations of features – such as object existence, object attributes, and object relations – during training, they develop a pronounced associative bias, which leads the models to expect these co-occurring features in similar situations. Consequently, when a model encounters an image containing only one element of a familiar combination, it may erroneously infer the presence of the associated feature. This associative bias is one primary source of model hallucinations Li et al. (2023b). To explore this phenomenon, we initially analyze the co-occurrence statistics of object-object, object-attribute, and object-relation-object combinations within the extensively annotated GQA Hudson and Manning (2019) dataset. We then curate a collection of images representing frequently and infrequently co-occur (object, object), (object, attribute), (object, relation, object) tuples. By doing so, we identify the most challenging images to construct a benchmark, to which we then add detailed human annotations for later thorough evaluation.
@@ 4 VALOR-Eval
We propose a framework VALOR-Eval that generalizes CHAIR, a metric that is widely adopted in existing studies Zhai et al. (2023); Wang et al. (2023b), by introducing semantic matching and incorporates both the faithfulness and coverage aspects into the evaluation. As shown in Figure 3, our evaluation process has two steps: feature extraction and matching (Section 4.1) and scoring (Section 4.2).
@@ 4.1 Feature Extraction and Matching
We start the process by generating an initial response, denoted as , using a specific LVLM with the input pair , where denotes the image and represents the prompt designed for LVLMs generation from VALOR-Bench. Following this, we leverage an LLM to analyze and extract key features. This is achieved through a series of prompts , outlined in Appendix D, which are designed to extract features from object existence, attributes, and relations, respectively, resulting in a comprehensive list of extracted features from , denoted as . Next, we utilize an LLM to align the extracted features list with a pre-annotated ground-truth features list from VALOR-Bench. This alignment is facilitated through a set of carefully crafted prompts , outlined in Appendix E, tailored to each feature subset, aiming to identify correlations and correspondences. Unlike previous evaluation metrics that rely on a fixed feature list and direct mapping, our approach eschews pre-processing and instead utilizes LLMs’ nuanced language comprehension capabilities to semantically match extracted features with their ground-truth counterparts. This process yields two key outputs: matched features dictionary () and broader conceptual matches dictionary ().
@@ 5.1 Model Coverage-Faithfulness Evaluation
In the evaluation of various models, as shown in Table 3, Emu2 distinguishes itself by achieving the highest average faithfulness score of 87.5, signifying its consistent capability to generate responses that accurately reflect the content of the input image. However, Emu2’s performance in terms of coverage is less impressive, with the lowest average score of 8.1, suggesting that its responses, while accurate, may not comprehensively cover all elements of the image. When broken down into specific dimensions, Emu2 excels in faithfulness across categories – scoring 94.2 in object existence, 54.3 in attribute-people, 72.2 in relation-positional, and 87.5 in relation-comparative. Conversely, it lags in coverage, with scores of 14.1 in object existence, 10.4 in attribute-object, 1.9 in attribute-people, and 1.8 in relation-positional. These results point to a potential trade-off between faithfulness and coverage in Emu2’s design, where the model prioritizes accuracy at the expense of a broader scope in its responses. This pattern supports the initial hypothesis that some LVLMs may intentionally sacrifice coverage to improve the precision of their outputs.
Meanwhile, GPT-4V distinguishes itself with an unparalleled average coverage score of 28.0, showcasing its adeptness in encapsulating a wide array of features from the input image. This indicates that GPT-4V excels in recognizing and addressing diverse elements within images, although it does not necessarily always maintain the highest accuracy, as seen in its lower faithfulness score of 61.6. Particularly in evaluations concerning the existence of objects, GPT-4V leads with the highest coverage score of 38.8, underlining its comprehensive approach to object detection. This approach tends to favor inclusivity, which might lead to the occasional identification of objects that are not present in the image. Furthermore, in evaluations focused on attributes related to people, GPT-4V again achieves the highest coverage score. However, this comes with a trade-off, as it also exhibits a higher tendency towards hallucinations compared to other models, indicating a propensity to generate details or elements that may not be grounded in the actual content of the image.
@@ 1. How does our co-occurrence data selection method compare to other alternatives?
To illustrate the effectiveness of the co-occurrence data selection method, we set up a baseline of randomly selecting 50 images in the GQA validation split and apply human annotations, the same as for our dataset. For the ablation study, we focus on the well-studied object hallucination. We evaluate three popular models representing query tokens-based image features (InstructBLIP), linear projection-based features (LLaVA-1.5), and advanced commercial LVLMs (GPT-4V). As shown in Table 5, all models tend to produce more hallucinations and exhibit significantly lower faithfulness compared to our benchmark. Notably, LLaVA-1.5 scores 12.4 points lower in faithfulness when evaluated against our benchmark.This suggests that our benchmark is challenging due to its reliance on co-occurrence selection. Additionally, the coverage scores for both LLaVA-1.5 and GPT-4V decreased. Upon further analysis through human review, we discover that our benchmark, on average, contains 1.69 more objects than images selected at random. This finding indicates that our data selection method can incorporate more complex objects compared to the random selection approach commonly used in other benchmark constructions.
@@ 7 Ethical Considerations
Our work investigates the phenomenon of hallucinations in outputs generated by LVLMs. Here, we outline the primary ethical considerations associated with our study. In developing our evaluation framework, we employed GPT-4 for feature extraction and matching tasks to evaluate the model’s hallucination. Consequently, we recognize that any biases inherent to the GPT-4 model will likely influence the results observed in our benchmark OpenAI (2023); Huang et al. (2023a); Qiu et al. (2023). Furthermore, our data collection efforts encompassed datasets from GQA and images sourced from the internet (specifically Pixel555https://www.pexels.com/). We acknowledge and adhere to the pertinent policies and requirements governing data sharing and utilization within our benchmark.
@@ Appendix A Large Vision-Language Models
Model | Visual Encoder | Alignment Network | Language Model |
InstructBLIP | Q-Former | ||
LLaVA-1.5 | MLP | ||
MiniGPT-v2 | Linear Projection | ||
mPLUG-Owl2 | Cross Attention | ||
BLIVA | Q-Former & Linear Projection | ||
CogVLM | MLP | ||
InternLM-Xcomposer2 | Partial Low-Rank Adaptation | ||
Qwen-VL | Cross Attention | ||
Emu2 | Linear Projection | ||
GPT-4(V) | Unknown | Unknown | GPT-4 |

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) suffer from hallucination issues, wherein the models generate plausible-sounding but factually incorrect outputs, undermining their reliability. A comprehensive quantitative evaluation is necessary to identify and understand the extent of hallucinations in these models. However, existing benchmarks are often limited in scope, focusing mainly on object hallucinations. Furthermore, current evaluation methods struggle to effectively address the subtle semantic distinctions between model outputs and reference data, as well as the balance between hallucination and informativeness. To address these issues, we introduce a multi-dimensional benchmark covering objects, attributes, and relations, with challenging images selected based on associative biases. Moreover, we propose a large language model (LLM)-based two-stage evaluation framework that generalizes the popular CHAIR metric Rohrbach et al. (2018) and incorporates both faithfulness and coverage into the evaluation. Experiments on 10 established LVLMs demonstrate that our evaluation metric is more comprehensive and better correlated with humans than existing work when evaluating on our challenging human-annotated benchmark dataset. Our work also highlights the critical balance between faithfulness and coverage of model outputs, and encourages future works to address hallucinations in LVLMs while keeping their outputs informative.111Our dataset and code can be found here: https://github.com/haoyiq114/VALOR.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Liu et al. (2023b); OpenAI (2023); Chen et al. (2023) have shown remarkable performance across a broad range of vision-language tasks. Despite the promising progress, the issue of hallucinations has emerged as a critical concern. Hallucination refers to the generation of plausible-sounding but inaccurate or fabricated textual descriptions for a given image, which can compromise the reliability and trustworthiness of the models.
@@
Recent studies have proposed various methods to evaluate models’ generative hallucinations Wang et al. (2023a); Zhai et al. (2023); Jing et al. (2023) and discriminative hallucinations Li et al. (2023b); Guan et al. (2023); Lovenia et al. (2023). However, they predominantly focus on hallucinations concerning object existence and their faithfulness within generated content, often neglecting other critical types of hallucinations and the assessment of coverage. This oversight can result in a lack of attention to the variety and depth of hallucinations that may occur beyond object identification, such as attributes and relations. Furthermore, these evaluation methods are often constrained by a predefined vocabulary, thus are inherently limited to fully appreciating the richness of the free-form generated captions. Specifically, the evaluation metrics may not capture novel expressions that extend beyond the predetermined vocabulary.
In contrast to prior studies, we introduce a human-annotated multi-dimensional evaluation benchmark VALOR-Bench222VALOR is short for vision-language attribute, relation, and object coverage and faithfulness. by breaking down hallucinations into three categories: object (existence), attributes (color and count), and relations (positional and comparative). In addition, to make the test cases challenging, we utilize the associative biases Li et al. (2023b); Zhou et al. (2023) presented in training datasets to select images with only one component of commonly co-occurred pairs or groups, leading models to mistakenly generate associated elements that are not present. Our experimental findings validate the effectiveness of this methodology in exposing the susceptibility of current LVLMs to such biases.
@@
We evaluate 10 mainstream LVLMs on VALOR-Bench, focusing on the balance between faithfulness and coverage score. We notice that even GPT-4V(ision) OpenAI (2023) still suffers from hallucination, achieving a relatively low faithfulness score despite covering more information within an image compared to other models.
@@ 2 Existing LVLMs Hallucination Evaluation Benchmarks and Metrics
As shown in Table 1, existing studies Li et al. (2023b); Wang et al. (2023a); Zhai et al. (2023); Lovenia et al. (2023); Villa et al. (2023); Petryk et al. (2024); Kaul et al. (2024) have primarily focused on object-level hallucination, with only a few recent studies Jing et al. (2023); Wang et al. (2023b); Jiang et al. (2024); Zhang et al. (2024) recognizing the importance of extending hallucinations to other dimensions. Our benchmark VALOR-Bench covers hallucination evaluations of objects, attributes, and relations, and we further detail attributes to color and counting, and relations to positional and comparative, to provide a comprehensive and fine-grained evaluation benchmark.
@@ 3.1 Image Collection
We aim to select images that can effectively expose the issue of model hallucinations. We hypothesize that when models are repeatedly exposed to specific combinations of features – such as object existence, object attributes, and object relations – during training, they develop a pronounced associative bias, which leads the models to expect these co-occurring features in similar situations. Consequently, when a model encounters an image containing only one element of a familiar combination, it may erroneously infer the presence of the associated feature. This associative bias is one primary source of model hallucinations Li et al. (2023b); Zhou et al. (2023). To explore this phenomenon, we initially analyze the co-occurrence statistics of object-object, object-attribute, and object-relation-object combinations within the extensively annotated GQA Hudson and Manning (2019) dataset. We then curate a collection of images representing frequently and infrequently co-occur (object, object), (object, attribute), (object, relation, object) tuples. By doing so, we identify the most challenging images to construct a benchmark, to which we then add detailed human annotations for later thorough evaluation.
@@ 4 VALOR-Eval
We propose a framework VALOR-Eval that generalizes CHAIR, a metric that is widely adopted in existing studies Zhai et al. (2023); Wang et al. (2023b), by introducing semantic matching and incorporating both the faithfulness and coverage aspects into the evaluation. As shown in Figure 3, our evaluation process has two steps: feature extraction and matching (Section 4.1) and scoring (Section 4.2).
@@ 4.1 Feature Extraction and Matching
We start the process by generating an initial response, denoted as , using a specific LVLM with the input pair , where denotes the image and represents the prompt designed for LVLMs generation from VALOR-Bench. Then, we leverage an LLM to analyze and extract key features. This is achieved through a series of prompts , outlined in Appendix D, which are designed to extract features from object existence, attributes, and relations, respectively, resulting in a comprehensive list of extracted features from , denoted as . Next, we utilize an LLM to align the extracted features list with a pre-annotated ground-truth features list from VALOR-Bench. This alignment is facilitated through a set of carefully crafted prompts , outlined in Appendix E, tailored to each feature subset, aiming to identify correlations and correspondences. Unlike previous evaluation metrics that rely on a fixed feature list and direct mapping, our approach eschews pre-processing and instead utilizes LLMs’ language comprehension capabilities to semantically match extracted features with their ground-truth counterparts. This process yields two key outputs: matched features dictionary () and broader conceptual matches dictionary ().
@@ 5.1 Model Coverage-Faithfulness Evaluation
In the evaluation of various models, as shown in Table 3, Emu2 distinguishes itself by achieving the highest average faithfulness score of 74.98, signifying its consistent capability to generate responses that accurately reflect the content of the input image. However, Emu2’s performance in terms of coverage is less impressive, with the lowest average score of 8.1, suggesting that its responses, while accurate, may not comprehensively cover all elements of the image. When broken down into specific dimensions, Emu2 excels in faithfulness across categories – scoring 94.2 in object existence, 54.3 in attribute-people, 72.2 in relation-positional, and 87.5 in relation-comparative. Conversely, it lags in coverage, with scores of 14.1 in object existence, 10.4 in attribute-object, 1.9 in attribute-people, and 1.8 in relation-positional. These results point to a potential trade-off between faithfulness and coverage in Emu2’s design, where the model prioritizes accuracy at the expense of a broader scope in its responses. This pattern supports the initial hypothesis that some LVLMs may intentionally sacrifice coverage to improve the precision of their outputs.
Meanwhile, GPT-4V(ision) distinguishes itself with an unparalleled average coverage score of 28.0, showcasing its adeptness in encapsulating a wide array of features from the input image. This indicates that GPT-4V excels in recognizing and addressing diverse elements within images, although it does not necessarily always maintain the highest accuracy, as seen in its lower faithfulness score of 61.6. Particularly in evaluations concerning the existence of objects, GPT-4V leads with the highest coverage score of 38.8, underlining its comprehensive approach to object detection. This approach tends to favor inclusivity, which might lead to the occasional identification of objects that are not present in the image. Furthermore, in evaluations focused on attributes related to people, GPT-4V again achieves the highest coverage score of 54.3. However, this comes with a trade-off, as it also exhibits a higher tendency towards hallucinations compared to other models, indicating a propensity to generate details or elements that may not be grounded in the actual content of the image.
@@ 1. How does our co-occurrence data selection method compare to other alternatives?
To illustrate the effectiveness of the co-occurrence data selection method, we set up a baseline of randomly selecting 50 images in the GQA validation split and applying human annotations, the same as for our dataset. For the ablation study, we focus on the well-studied object hallucination. We evaluate three popular models representing query tokens-based image features (InstructBLIP), linear projection-based features (LLaVA-1.5), and advanced commercial LVLMs (GPT-4V). As shown in Table 5, all models tend to produce more hallucinations and exhibit significantly lower faithfulness compared to our benchmark. Notably, LLaVA-1.5 scores 12.4 points lower in faithfulness when evaluated against our benchmark. This suggests that our benchmark is challenging due to its reliance on co-occurrence selection. Additionally, the coverage scores for both LLaVA-1.5 and GPT-4V decreased. Upon further analysis through human review, we discover that our benchmark, on average, contains 1.69 more objects than images selected at random. This finding indicates that our data selection method can incorporate more complex objects compared to the random selection approach commonly used in other benchmark constructions.
@@ 7 Ethical Considerations
Our work investigates the phenomenon of hallucinations in outputs generated by LVLMs. Here, we outline the primary ethical considerations associated with our study. In developing our evaluation framework, we employed GPT-4 for feature extraction and matching tasks to evaluate the model’s hallucination. Consequently, we recognize that any biases inherent to the GPT-4 model will likely influence the results observed in our benchmark OpenAI (2023); Huang et al. (2023a); Qiu et al. (2023); Wang et al. (2024b). Furthermore, our data collection efforts encompassed datasets from GQA and images sourced from the internet (specifically Pixel555https://www.pexels.com/). We acknowledge and adhere to the pertinent policies and requirements governing data sharing and utilization within our benchmark.
@@ 8 Limitations
9 Acknowledgment
We thank anonymous reviewers for their helpful feedback. We also thank members from the UCLA NLP group for their feedback and discussions. This research is supported by Meta Sponsor Research Award, Okawa Foundation Research Grant, Google Research Scholar, and a gift from UCLA Institute for Technology, Law and Policy. This material is based on research supported by the ECOLE program under Cooperative Agreement HR00112390060, both with the US Defense Advanced Research Projects Agency (DARPA).